OpenStudy (anonymous):

A point charge q1=4.30 nC is located on the x-axis at x=2.30m, and a second point charge q2=-6.00 nC is on the y-axis at y=1.15m. (a) What is the total electric flux due to these two point charges through a spherical surface centered at the origin and with radius r1=0.615m? (b) with radius r2=1.55m? (c) with radius r3=2.45m? I really honestly only need one explained to me, after I see it worked out once, I'm pretty sure I can do it again.

OpenStudy (anonymous):

Okay,I've figured out that a sphere r1 is not affected by these charges because it does not enclose either of them. r2 eencloses only q2, and r3 encloses q1 and q2. Therefore: (a) Φ =0 (b) Φ= q2/Eo =(-6.00*10^-9)/(8.854*10^-12) =-677.659 N*m^2/C (c) Φ= (q1+q2)/Eo =[(4.30*10^-9)+(-6.00*10^-9)]/(8.854*10^-12) =-192.004 N*m^2/C